OpenStudy (anonymous):

Find the sum of the first 38 terms of the arithmetic sequence: 3, 8, 13, 18, 23, ... A) Sn= 988 B) Sn= 1010 C) Sn= 3283 D) Sn= 3629

OpenStudy (anonymous):

help

OpenStudy (callisto):

it is a AS with d=5 a=3 therefore the general term = a +d(n-1) = 3+5(n-1) = 5n-2 the 38term = 5(38)-2 = 188 sum of first 38 terms = (a+l)n/2 = (3+188)x38/2 = 3629